# MOTGNN: Interpretable Graph Neural Networks for Multi-Omics Disease Classification

### v1

@@ Abstract.
Integrating multi-omics data, such as DNA methylation, mRNA expression, and microRNA (miRNA) expression, offers a comprehensive view of the biological mechanisms underlying disease. However, the high dimensionality and complex interactions among omics layers present major challenges for predictive modeling. We propose Multi-Omics integration with Tree-generated Graph Neural Network (MOTGNN), a novel and interpretable framework for binary disease classification. MOTGNN employs eXtreme Gradient Boosting (XGBoost) to perform omics-specific supervised graph construction, followed by modality-specific Graph Neural Networks (GNNs) for hierarchical representation learning, and a deep feedforward network for cross-omics integration. On three real-world disease datasets, MOTGNN outperforms state-of-the-art baselines by 5-10% in accuracy, ROC-AUC, and F1-score, and remains robust to severe class imbalance (e.g., 87.2% vs. 33.4% F1 on imbalanced data). The model maintains computational efficiency through sparse graphs (2.1-2.8 edges per node) and provides built-in interpretability, revealing both top-ranked biomarkers and the relative contributions of each omics modality. These results highlight MOTGNN’s potential to improve both predictive accuracy and interpretability in multi-omics disease modeling.
@@ 1. Introduction
A comprehensive understanding of complex diseases, such as cancer, cardiovascular disease, and neurodegenerative disorders, is critical for advancing public health and enabling precision medicine. These diseases are not driven by isolated molecular events but emerge from intricate interactions among genomic, epigenomic, and transcriptomic factors (Subramanian et al., 2020). Recent advances in high-throughput biotechnologies have enabled the simultaneous collection of different layers of molecular regulation, such as gene expression (messenger RNA), epigenetic modifications (DNA methylation), and non-coding RNA activity (e.g., microRNA), from the same set of biological samples (Hood and Rowen, 2013). These omics datasets capture intricate relationships among biological entities and molecular processes, offering insights into disease mechanisms. While individual omics types provide valuable information, different modalities reflect distinct biological processes and contribute complementary information. Integrating multiple omics (multi-omics) data is essential to comprehensively understand the complex, multi-layered systems underlying disease processes and their progression (Waqas et al., 2024). However, integrating heterogeneous omics data presents unique challenges. A fundamental issue is that the number of samples ($n$) is substantially smaller than the number of features ($p$). For example, a study may involve only a few hundred participants while measuring tens of thousands of molecular features. This high dimensionality complicates learning and increases the risk of overfitting. Moreover, not all measured features are biologically relevant; noisy or redundant features can obscure meaningful patterns. Although prior biological knowledge, such as gene-gene or protein-protein interactions, has been identified experimentally, our knowledge is constrained by experimental design and equipment constraints. Consequently, there is a growing need for computational models that can reliably infer biologically meaningful structures and interactions from multi-omics data (Jiang et al., 2025).
@@
Traditional machine learning and deep learning models have been extensively applied to address the high dimensionality and complexity of omics data (Ayman et al., 2023; Zaghlool and Attallah, 2022), but most of these approaches assume a Euclidean data structure. Classical machine learning methods, such as random forests (Breiman, 2001) and gradient boosting methods, such as XGBoost (Chen and Guestrin, 2016), operate by constructing ensembles of decision trees and are effective for classification tasks, but they typically treat features as independent and fail to capture structural relationships among them. Deep learning models (Lecun et al., 2015), such as deep feedforward networks (DFNs) (Goodfellow et al., 2016), convolutional neural networks, and recurrent neural networks, can model nonlinear feature dependencies but are not designed to process non-Euclidean data structures such as graphs. These approaches typically require large training sets to avoid overfitting. Moreover, their black-box nature often limits interpretability, which is essential in biomedical applications where transparency and reproducibility are critical. Furthermore, concatenating features from different omics modalities as input can obscure modality-specific relationships and introduce additional noise. Graph-based approaches offer a natural alternative for modeling complex biological systems. In omics applications, nodes can represent features (e.g., genes) or samples, and edges encode various relationships, such as similarity, co-expression, or learned interactions. Graph Neural Networks (GNNs) (Bronstein et al., 2017; Wu et al., 2020; Zhang et al., 2020; Zhou et al., 2020; Chen et al., 2023) have emerged as powerful tools for learning from graph-structured data by recursively aggregating information from each node’s neighborhood. GNNs capture both local and global graph topology and have found increasing use in biomedical applications, including molecular modeling, medical imaging, drug discovery, and disease prediction (Zhang et al., 2021; Johnson et al., 2024; Waqas et al., 2024; Zhong et al., 2025). Foundational models such as message passing neural networks (MPNNs) (Gilmer et al., 2017), graph convolutional networks (GCNs) (Kipf and Welling, 2017), graph attention networks (GATs) (Petar et al., 2018), GraphSAGE (Hamilton et al., 2017), and graph isomorphism networks (GINs) (Xu et al., 2019) have significantly advanced the expressiveness and scalability of GNN-based learning.
Despite recent advances, applying GNNs to multi-omics data remains challenging and relatively underexplored (Zhang et al., 2021; Valous et al., 2024; Paul et al., 2024; Ballard et al., 2024). The irregular nature of graph-structured data further complicates this task, as each node may have a variable number of neighbors, and many standard deep learning operations are not directly applicable in this setting (Chen et al., 2023). Although several GNN-based frameworks have been proposed for multi-omics integration, many of these works exhibit substantial limitations, which we summarize into three main areas: (1) Inadequate modeling of high-dimensional, heterogeneous omics data: To address high dimensionality (i.e., the “small $n$, large pp” issue), existing methods typically apply simple filtering techniques, such as removing low-variance or zero-valued features, or selecting features based on univariate statistical tests. Some approaches flatten multi-omics data by concatenating omics layers before graph construction, which weakens modality-specific signals and ignores biological heterogeneity. Others rely heavily on hand-crafted similarity graphs (e.g., Pearson or cosine similarity), which may miss nonlinear, task-specific relationships, ignore the sparse and scale-free nature of biological graphs, and are sensitive to arbitrary thresholding or preprocessing pipelines. (2) Vulnerability to class imbalance: While class imbalance is prevalent in biomedical datasets (e.g., rare diseases, early-stage cancer detections, or minority subtypes), it is rarely addressed during model development or evaluation. Many existing models optimize for overall accuracy, which can be misleading when predictions favor the majority class. In these scenarios, even high accuracy may reflect poor minority class detection. As a result, these approaches often lack robustness and generalizability in imbalanced settings, where accurate prediction of minority cases is of greater biological and clinical importance. (3) Limited interpretability and biological insight: Interpretability is essential in biomedical research, as understanding which features or omics modalities drive predictions is essential for trust, transparency, and downstream biological discovery. However, most existing GNN-based models lack built-in interpretability. Instead, they rely on computationally intensive post hoc analyses or ablation studies to infer importance, which are not scalable and may introduce instability. In addition, they typically treat all omics types equally, failing to uncover their relative contributions. To address these limitations, we propose MOTGNN, a novel and interpretable graph-based framework for multi-omics data integration that combines supervised graph construction with hierarchical representation learning. Our key contributions are as follows:
  • Modality-specific modeling with supervised graph learning: We utilize XGBoost to filter informative features for each omics type, and leverage the structure of the trained decision trees to construct sparse, supervised, and modality-specific graphs. This approach preserves biologically meaningful interactions while reducing noise and redundancy. We then apply GNNs to learn latent representations from each graph, which are integrated through a DFN to model cross-omics interactions.
  • Imbalance-robust architecture: MOTGNN consistently outperforms existing models across all datasets, achieving 10-50% improvement in F1-score on class-imbalanced real-world data. Its design mitigates overfitting to the dominant classes and remains effective in detecting minority class signals, addressing a critical limitation in biomedical classification tasks.
  • Integrated interpretability at feature and omics levels: MOTGNN enables end-to-end interpretability without requiring post hoc analysis. It provides (i) feature-level importance scores to identify top-ranked biomarkers within each modality, and (ii) omics-level contribution scores, offering insights into which data types (e.g., methylation, mRNA, or miRNA) are most influential for prediction. This layered interpretability potentially supports downstream biological discovery and facilitates clinical relevance.
Figure 1. Overview of the proposed MOTGNN framework, which involves three key modules: (i) XGBoost for graph construction within each omics modality; (ii) GNN for learning modality-specific embeddings by encoding both feature graphs and associated data matrices; and (iii) DFN that integrates the learned embeddings and performs classification.
[…]
[MOTGNN architecture]The MOTGNN framework includes XGBoost for graph construction, a GNN module for encoding, and a DFN for integration and classification.
2. Related Work
Many advanced deep learning methods have been developed by incorporating novel model components to analyze omics data. For example, MOCAT (Yao et al., 2024) leverages an autoencoder, three omics-specific classifiers, a multi-head attention mechanism, and a confidence network (ConfNet) to integrate and learn from multi-omics data (e.g., methylation, mRNA, and miRNA). However, it lacks the ability to capture topological information and thus may miss important structural insights. Some GNN-based methods have been developed for single-omics data analysis in disease classification or cancer subtype prediction. For example, forgenet (Kong and Yu, 2020) constructs a supervised directed graph using a tree-ensemble model, but it uses the original high-dimensional data without dimension reduction, leading to high computational costs. Moreover, it only uses a single omics type, which provides limited information compared to multi-omics integration. DMGCN (Yang et al., 2021) integrates multiple biological networks via GCNs, but also focuses on a single omics modality and does not utilize cross-omics information.
@@ 2. Related Work
More recent GNN-based methods aim to integrate multi-omics data for disease classification, cancer subtype analysis, and survival prediction. MOGONET (Wang et al., 2021) proposes a GCN-based framework to integrate multi-omics data (methylation, mRNA, and miRNA), using a cross-omics discovery tensor and a view correlation discovery network for final label prediction. However, it identifies important features through computationally expensive ablation studies. In addition, its graph construction is based on cosine similarity with a manually selected threshold determined by a hyperparameter (edge-to-node ratio), which must be tuned. MODILM (Zhong et al., 2023) shares a similar architecture with MOGONET but uses GAT instead of GCN. CLCLSA (Zhao et al., 2024) handles incomplete multi-omics classification using cross-omics autoencoders to recover missing omics representations, contrastive learning to align representations, and self-attention at both the feature and omics levels for feature selection, but does not explicitly leverage graph structures. SUPREME (Kesimoglu and Bozdag, 2023) employs GCNs and patient similarity networks, integrating seven types of omics data for cancer subtype prediction and survival analysis. However, it also relies on ablation studies to determine the most effective omics combinations. MOGAT (Tanvir et al., 2024) focuses on survival analysis with a similar structure to SUPREME, but uses GAT instead of GCN. It conducted integration studies with two to eight omics types and found that using three omics achieved the best performance for breast cancer subtype prediction, suggesting that adding more omics data does not always yield better results. However, MOGAT does not address model interpretability, as it does not identify important biomarkers. DeepMoIC (Wu et al., 2024) integrates three types of omics data to address cancer subtype classification by combining autoencoders, similarity network fusion, and a deep GCN with residual connections and identity mappings. However, it constructs patient similarity networks to guide the learning and lacks omics-wise interpretability. MoGCN (Li et al., 2022) targets breast cancer subtype classification using multi-omics data (copy number variation, RPPA, and RNA-seq). It applies autoencoders for dimensionality reduction and similarity network fusion for graph construction, and trains a GCN model on the resulting fused network. However, its graph construction still relies on distance-based methods.
While these models demonstrate promising results, several common limitations remain: They require extensive data preprocessing, and the quality of such preprocessing can significantly affect downstream model performance. Their graph construction approaches rely on simple similarity metrics, which often fail to capture task-relevant nonlinear interactions and can introduce bias due to arbitrary thresholding. They rarely address the class imbalance issue, which is prevalent in biomedical datasets and can significantly impact performance. They commonly depend on ablation studies to identify the optimal combination of omics data types, a computationally expensive process that is not easily scalable.
@@ 3. Methodology
We summarized the key notations and their associated descriptions in Table 1. A graph is denoted as $G(V,E)$, where $V$ represents the set of vertices (or nodes) and $E$ denotes the set of edges. Graph structure is encoded by its adjacency matrix $A$, in which the $(i,j)$-th element is one if node $i$ is connected to node $j$, and zero otherwise. For undirected graphs, $A$ is symmetric. The adjacency matrix augmented with self-loops is denoted as $\tilde{A}$, where all diagonal elements are set to one to include self-connections.
Table 1. Key notations and descriptions
@@ 3.1. Proposed Model
We propose a Multi-Omics integration framework with Tree-generated Graph Neural Network (MOTGNN) model for binary disease classification. MOTGNN first leverages eXtreme Gradient Boosting (XGBoost) to select informative features in a supervised manner, as the tree-based feature selection captures strong predictive signals. It then utilizes GNNs to learn omics-specific representations from graphs constructed using the structure of trained XGBoost trees. This combination allows the model to extract both local and global patterns from each omics layer while maintaining biological interpretability.
The framework of MOTGNN is shown in Figure 1, which consists of three primary modules: (1) An XGBoost module for feature graph construction; (2) A GNN module to encode each omics-specific graph and its associated data matrix; and (3) A DFN module to integrate and classify the learned embeddings. This modular architecture allows MOTGNN to handle high-dimensional, heterogeneous multi-omics data efficiently while providing interpretable insights at both the feature and omics-layer levels. Algorithm 1 demonstrates the steps of the overall structure of MOTGNN.
@@
9:Concatenate $Z_{1}$, $Z_{2}$, and $Z_{3}$ to form a unified representation
10:Pass the concatenated representation $Z$ into a DFN to perform final binary classification
@@ 3.2. Graph Construction
XGBoost (Chen and Guestrin, 2016) is a powerful ensemble method based on gradient-boosted decision trees. Compared to traditional gradient boosting machines (GBMs) (Friedman, 2002), XGBoost introduces a regularized learning objective to reduce overfitting and improve generalization. In MOTGNN, we employ three XGBoost models, each trained independently to perform binary classification on the omics-specific datasets $X_{1}$, $X_{2}$, and $X_{3}$ alongside the binary response $Y$. Unlike random forest, which generates independent trees, each trained XGBoost model generates $M$ sequential decision trees, where every tree improves upon the residuals of the previous ones. During training, each tree selects a subset of features for node splits. We collect the union of all features used across the $M$ trees and define these as the selected features. This yields a supervised, tree-guided feature selection process that reduces the original feature dimensions from $p_{1}$, $p_{2}$, and $p_{3}$ to $p_{1}^{*}$, $p_{2}^{*}$, and $p_{3}^{*}$, respectively, where $p_{i}^{*}<p_{i}$ and $p_{1}^{*}\neq p_{2}^{*}\neq p_{3}^{*}$. The reduced dimensions preserve features with higher predictive importance, facilitating downstream graph construction and neural modeling.
Inspired by forgeNet (Kong and Yu, 2020), which constructs feature graphs based on the structure of ensemble trees, we adapt this approach to build omics-specific undirected graphs from XGBoost models. Specifically, for each omics dataset, we treat each trained decision tree as a graph, where features involved in splits are nodes, and parent-child relationships form the undirected edges. Taking $X_{1}\in\mathbb{R}^{n\times p_{1}}$ along with the binary outcome $Y$ as an example, suppose the XGBoost model generates $M$ sequential decision trees. Each tree is considered an undirected graph $G^{m}=(V^{m},E^{m})$ for $m=1,2,\ldots,M$. We then take the union of all tree-level graphs to form the aggregated graph $G_{1}=(V_{1},E_{1})$ as:
@@
where the number of nodes in $G_{1}$ is $|V_{1}|=p_{1}^{*}$, and $p_{1}^{*}<p_{1}$. We also add self-loops to the adjacency matrices to preserve self-information in GNN message passing. A similar procedure generates $G_{2}$ and $G_{3}$ for the remaining two omics datasets.

### v2

@@ Abstract.
Integrating multi-omics data, such as DNA methylation, mRNA expression, and microRNA (miRNA) expression, offers a comprehensive view of the biological mechanisms underlying disease. However, the high dimensionality of multi-omics data, the heterogeneity across modalities, and the lack of reliable biological interaction networks make meaningful integration challenging. In addition, many existing models rely on handcrafted similarity graphs, are vulnerable to class imbalance, and often lack built-in interpretability, limiting their usefulness in biomedical applications. We propose Multi-Omics integration with Tree-generated Graph Neural Network (MOTGNN), a novel and interpretable framework for binary disease classification. MOTGNN employs eXtreme Gradient Boosting (XGBoost) for omics-specific supervised graph construction, followed by modality-specific Graph Neural Networks (GNNs) for hierarchical representation learning, and a deep feedforward network for cross-omics integration. Across three real-world disease datasets, MOTGNN outperforms state-of-the-art baselines by 5-10% in accuracy, ROC-AUC, and F1-score, and remains robust to severe class imbalance. The model maintains computational efficiency through the use of sparse graphs and provides built-in interpretability, revealing both top-ranked biomarkers and the relative contributions of each omics modality. These results highlight the potential of MOTGNN to improve both predictive accuracy and interpretability in multi-omics disease modeling.
@@ 1. Introduction
A comprehensive understanding of complex diseases, such as cancer, cardiovascular disease, and neurodegenerative disorders, is critical for advancing public health and enabling precision medicine. These diseases are not driven by isolated molecular events but instead emerge from intricate interactions among genomic, epigenomic, and transcriptomic factors (Subramanian et al., 2020). Recent advances in high-throughput biotechnologies have enabled the simultaneous collection of multiple molecular layers, such as gene expression (mRNA), epigenetic modifications (DNA methylation), and non-coding RNA activity (e.g., microRNA), from the same set of biological samples (Hood and Rowen, 2013). These omics datasets capture intricate relationships among biological entities and molecular processes, offering insights into disease mechanisms. While individual omics types provide valuable information, each reflects distinct biological processes and contributes complementary information. Integrating multiple omics (multi-omics) data is therefore essential for comprehensively understanding the multi-layered systems underlying disease processes and progression (Waqas et al., 2024). However, integrating heterogeneous omics data presents unique challenges. A fundamental issue is that the number of samples ($n$) is typically much smaller than the number of features ($p$). For example, a study may involve only a few hundred participants while measuring tens of thousands of molecular variables. This high dimensionality complicates model training and increases the risk of overfitting. Moreover, not all measured features are biologically relevant; noisy or redundant features can obscure meaningful patterns. Although biological interaction networks such as gene-gene or protein-protein interactions have been identified experimentally, current knowledge is constrained by limitations in experimental design and equipment. Consequently, there is a growing need for computational models that can reliably infer biologically meaningful structures and interactions directly from multi-omics data (Jiang et al., 2025).
@@
Traditional machine learning and deep learning models have been widely applied to address the high dimensionality and complexity of omics data (Zaghlool and Attallah, 2022; Ayman et al., 2023), but most assume a Euclidean data structure. Classical methods such as random forests (Breiman, 2001) and gradient boosting algorithms like XGBoost (Chen and Guestrin, 2016) construct ensembles of decision trees and perform well for classification, yet they typically treat features as independent and fail to capture structural relationships among them. Deep learning models (Lecun et al., 2015), including deep feedforward networks (DFNs) (Goodfellow et al., 2016), convolutional neural networks, and recurrent neural networks, can model nonlinear feature dependencies but are not designed to process non-Euclidean data structures such as graphs. These approaches typically require large training sets to prevent overfitting and are often criticized for their black-box nature, which limits interpretability, an essential requirement in biomedical research where transparency and reproducibility are critical. Furthermore, simply concatenating features from different omics modalities into a single input can obscure modality-specific relationships and introduce additional noise. Graph-based approaches provide a natural alternative for modeling complex biological systems. In omics applications, nodes can represent features (e.g., genes) or samples, and edges encode various relationships such as similarity, co-expression, or learned interactions. Graph Neural Networks (GNNs) (Bronstein et al., 2017; Wu et al., 2020; Zhang et al., 2020; Zhou et al., 2020) have emerged as powerful tools for learning from graph-structured data by recursively aggregating information from each node’s neighborhood. GNNs capture both local and global graph topological patterns and have shown increasing success in biomedical domains, including molecular modeling, drug discovery, and disease prediction (Zhang et al., 2021; Johnson et al., 2024; Waqas et al., 2024; Zhong et al., 2025). Foundational architectures such as message passing neural networks (MPNNs) (Gilmer et al., 2017), graph convolutional networks (GCNs) (Kipf and Welling, 2017), graph attention networks (GATs) (Petar et al., 2018), GraphSAGE (Hamilton et al., 2017), and graph isomorphism networks (GINs) (Xu et al., 2019) have advanced the expressiveness and scalability of GNN-based learning.
@@
Recent GNN-based methods have aimed to integrate multi-omics data for disease classification, cancer subtype analysis, and survival prediction. MOGONET (Wang et al., 2021) introduces a GCN-based framework for integrating methylation, mRNA, and miRNA through a cross-omics discovery tensor and a view correlation discovery network. However, it identifies important features through computationally intensive ablation studies, and its graph construction relies on cosine similarity with a manually tuned edge-to-node ratio. MODILM (Zhong et al., 2023) adopts a similar structure but replaces GCN with GAT, while CLCLSA (Zhao et al., 2024) addresses incomplete omics data using cross-omics autoencoders, contrastive learning, and attention mechanisms without explicitly modeling graph structures. SUPREME (Kesimoglu and Bozdag, 2023) and MOGAT (Tanvir et al., 2024) extend graph-based modeling to cancer subtype prediction and survival analysis by employing patient similarity networks derived from predefined metrics. DeepMoIC (Wu et al., 2024) integrates three omics types using autoencoders, similarity network fusion, and a deep GCN with residual connections, while MoGCN (Li et al., 2022) focuses on breast cancer subtyping with graphs constructed via similarity network fusion. Although these models achieve promising results within their respective applications, they primarily rely on unsupervised or distance-based graphs and often lack built-in interpretability or robustness under class imbalance.
Despite recent advances, applying GNNs to multi-omics data remains challenging and relatively underexplored (Zhang et al., 2021; Valous et al., 2024; Paul et al., 2024; Ballard et al., 2024). Existing methods exhibit three key limitations: (1) Inadequate modeling of high-dimensional, heterogeneous omics data: To address the “small $n$, large pp” issue, many methods apply simple filtering (e.g., removing low-variance or zero-valued features) or flatten multi-omics data through concatenation, which weakens modality-specific signals and ignores biological heterogeneity. Others rely on handcrafted similarity graphs (e.g., Pearson or cosine), which may miss nonlinear, task-specific relationships and are sensitive to arbitrary thresholding. (2) Vulnerability to class imbalance: While class imbalance is common in biomedical datasets (e.g., rare diseases, early-stage cancer detections, or minority subtypes), it is rarely addressed during model development or evaluation. Many models emphasize overall accuracy, which can obscure poor minority-class performance and limit generalizability. (3) Limited interpretability and biological insight: Interpretability is essential in biomedical research for trust, transparency, and downstream discovery. However, most GNN-based models lack built-in interpretability and instead rely on computationally intensive post-hoc analyses or ablation studies that may be unstable. Furthermore, they typically treat all omics modalities equally, overlooking their relative contributions to prediction.
To address these limitations, we propose MOTGNN, a novel and interpretable graph-based framework for multi-omics data integration that combines supervised graph construction with hierarchical representation learning. Our main contributions are summarized as follows: (1) Modality-specific modeling with supervised graph learning: MOTGNN employs XGBoost to identify informative features for each omics type and uses the trained decision-tree structure to construct sparse, supervised, and modality-specific graphs. This approach preserves biologically meaningful interactions while reducing noise and redundancy. Omics-specific GNNs learn latent representations from each graph, which are then integrated through a deep feedforward network (DFN) to capture cross-omics interactions. (2) Imbalance-robust architecture: MOTGNN consistently outperforms existing models across all datasets, achieving 10-50% improvement in F1-score on class-imbalanced real-world data. Its design mitigates overfitting to dominant classes and effectively detects signals from minority classes, addressing a critical limitation in biomedical classification tasks. (3) Integrated interpretability at feature and omics levels: MOTGNN provides end-to-end interpretability without requiring post-hoc analysis. It generates (i) feature-level importance scores to identify top-ranked biomarkers within each modality, and (ii) omics-level contribution scores to quantify the relative importance of each data type (e.g., methylation, mRNA, or miRNA), supporting both biological discovery and clinical insight.
@@ 2. Methodology
We summarize the key notations and their associated descriptions in Table 1. A graph is denoted as $G(V,E)$, where $V$ represents the set of vertices (or nodes) and $E$ denotes the set of edges. The graph structure is encoded by its adjacency matrix $A$, where the $(i,j)$-th element is one if node $i$ is connected to node $j$, and zero otherwise. For undirected graphs, $A$ is symmetric. The adjacency matrix augmented with self-loops is denoted as $\tilde{A}$, where all diagonal elements are set to one to include self-connections.
Table 1. Key notations and their descriptions used throughout the paper.
@@
Figure 1. Overview of the proposed MOTGNN framework for multi-omics data integration and disease classification. The model comprises three key modules: (i) XGBoost for constructing omics-specific supervised feature graphs; (ii) graph neural network (GNN) for learning modality-specific embeddings by encoding each graph and its corresponding data matrix; and (iii) deep feedforward network (DFN) for integrating the learned embeddings and performing final classification.
@@ 2.1. Proposed Model
We propose MOTGNN, a Multi-Omics integration framework with Tree-generated Graph Neural Networks for binary disease classification. MOTGNN first leverages eXtreme Gradient Boosting (XGBoost) to select informative features in a supervised manner, as tree-based feature selection captures strong predictive signals. It then utilizes GNNs to learn omics-specific representations from graphs constructed using the structure of the trained XGBoost trees. This combination allows the model to extract both local and global patterns from each omics layer while maintaining biological interpretability.
The overall framework of MOTGNN (Figure 1) consists of three primary modules: (1) an XGBoost module for feature graph construction; (2) a GNN module to encode each omics-specific graph and its associated data matrix; and (3) a DFN module to integrate and classify the learned embeddings. This modular architecture allows MOTGNN to handle high-dimensional, heterogeneous multi-omics data efficiently, while providing interpretable insights at both the feature and omics levels. Algorithm 1 outlines the overall procedure of MOTGNN.
@@
9:Concatenate $Z_{1}$, $Z_{2}$, and $Z_{3}$ to form a unified representation $Z=[Z_{1}|Z_{2}|Z_{3}]$
10:Pass $Z$ into a DFN to perform final binary classification
@@ 2.2. Graph Construction
XGBoost (Chen and Guestrin, 2016) is a powerful ensemble method based on gradient-boosted decision trees. Compared with traditional gradient boosting machines, it introduces a regularized learning objective to reduce overfitting and improve generalization. In MOTGNN, we employ three XGBoost models, each trained independently to perform binary classification on the omics-specific datasets $X_{1}$, $X_{2}$, and $X_{3}$ together with the binary response $Y$. Unlike random forests, which generate independent trees, each XGBoost model generates $M$ sequential decision trees, where every tree improves upon the residuals of the previous ones. During training, each tree selects a subset of features for node splits. We collect the union of all features used across the $M$ trees and define these as the selected features. This yields a supervised, tree-guided feature selection process that reduces the original feature dimensions from $p_{1}$, $p_{2}$, and $p_{3}$ to $p_{1}^{*}$, $p_{2}^{*}$, and $p_{3}^{*}$, respectively, where $p_{i}^{*}<p_{i}$ and, in general, $p_{1}^{*}\neq p_{2}^{*}\neq p_{3}^{*}$. The reduced feature spaces preserve variables with higher predictive importance, facilitating downstream graph construction and neural modeling.
Inspired by forgeNet (Kong and Yu, 2020), which constructs feature graphs based on the structure of ensemble trees, we adapt this idea to build omics-specific undirected graphs from XGBoost models. For each omics dataset, we treat each trained decision tree as a graph, where features involved in splits are represented as nodes and parent-child relationships form undirected edges. For example, consider $X_{1}\in\mathbb{R}^{n\times p_{1}}$ and the binary outcome $Y$: if the XGBoost model generates $M$ sequential decision trees, each tree is viewed as an undirected graph $G^{m}=(V^{m},E^{m})$ for $m=1,2,\ldots,M$. We then take the union of all tree-level graphs to form an aggregated graph as
@@
where the number of nodes is $|V_{1}|=p_{1}^{*}$, and $p_{1}^{*}<p_{1}$. Self-loops are added to the adjacency matrices to preserve self-information in GNN message passing. The similar procedure is used to generate $G_{2}$ and $G_{3}$ for the remaining two omics datasets.